# Bayesian Inference Under Differential Privacy: Prior Selection Considerations with Application to Univariate Gaussian Data and Regression

### v1

@@ 5 Non-Univariate Gaussian Application
The strategies we discuss can be applied to any Bayesian analysis under DP. To demonstrate, we consider the DP Bayesian linear regression method of [5]. In particular, we adapt an example from [5] using data from [6] to illustrate how the strategies for enforcing constraints in Section 3 can be used off-the-shelf to supplement an existing method.
The data in [5, 6] include the response variable cirrhosis rate, $\{y_{i}\}$, explanatory variable drinking rate, $\{x_{i}\}$, and sample size $n=46$. The authors rescale both variables to lie in $[0,1]$ and assume the range of both quantities is public information. Letting ${\mathbf{Y}}\in{\mathbb{R}}^{n}$ be the vector of responses and ${\mathbf{X}}\in{\mathbb{R}}^{n\times 2}$ be a design matrix with a column of ones, the authors’ model is a simple linear regression of the form below with the following informative, conjugate prior.
${\mathbf{Y}}\sim{\mathcal{N}}_{n}({\mathbf{X}}{\boldsymbol{\theta}},\sigma^{2}{\mathbf{I}}_{n}),\qquad{\boldsymbol{\theta}}\mid\sigma^{2}\sim{\mathcal{N}}_{2}\left(\begin{bmatrix}1\\ 0\end{bmatrix},\frac{\sigma^{2}}{4}\mathbf{I}_{2}\right),\quad\sigma^{2}\sim\textsf{IG}\left(\frac{40}{2},\frac{1}{2}\right),$ (5)
where $\mathbf{I}_{p}$ is the $p\times p$ identity matrix. Their proposed method involves applying the Laplace Mechanism to each element of the sufficient statistics ${\mathbf{X}}^{\top}{\mathbf{X}}$, ${\mathbf{X}}^{\top}{\mathbf{Y}}$ and ${\mathbf{Y}}^{\top}{\mathbf{Y}}$. To ensure a valid regression solution exists, the authors enforce that the matrix ${\mathbf{Z}}^{\top}{\mathbf{Z}}$, where ${\mathbf{Z}}=[{\mathbf{X}}\hskip 5.78172pt{\mathbf{Y}}]$, is positive semi-definite (PSD) by projecting sampled values to the nearest PSD matrix. They apply their Gibbs-SS-Noisy method, which also adds Laplace noise to the matrix of fourth sample moments of ${\mathbf{X}}$.
Since it is known that $x_{i},y_{i}\in[0,1]$ for all $i$, we can exploit the resulting structure in the linear regression to constrain parameter updates in the Gibbs sampler. In particular, letting ${\mathbf{X}}=[{\mathbf{1}}_{n}\hskip 5.78172pt{\mathbf{x}}_{1}]$, we show in Appendix I that $0\leq{\mathbf{x}}_{1}^{\top}{\mathbf{x}}_{1}\leq{\mathbf{x}}_{1}^{\top}{\mathbf{1}}\leq n$ and $0\leq{\mathbf{Y}}^{\top}{\mathbf{Y}}\leq{\mathbf{Y}}^{\top}{\mathbf{1}}\leq n$. We also show that ${\mathbf{x}}_{1}^{\top}{\mathbf{Y}}\leq\min\{{\mathbf{x}}_{1}^{\top}{\mathbf{1}},{\mathbf{Y}}^{\top}{\mathbf{1}}\}$ and that $40\leq\sigma^{2}\leq 1/$ 4. Letting the regression coefficients be ${\boldsymbol{\theta}}=[\theta_{0}\hskip 5.78172pt\theta_{1}]^{\top}$, we know that $(\theta_{1},\theta_{0})$ must lie in the region depicted in the right panels of Figure 5. We create a constrained Gibbs sampler by drawing each iterate from the unconstrained full conditional distribution and rejecting and resampling whenever all constraints are not satisfied. We also reject and resample if ${\mathbf{Z}}^{\top}{\mathbf{Z}}$ is not PSD.
@@
Figure 5: Plot of posterior draws from the linear regression method of [5]. The left panels represent draws of the imputed sufficient statistics ${\mathbf{Y}}^{\top}{\mathbf{1}}$ and ${\mathbf{Y}}^{\top}{\mathbf{Y}}$, while the right panels represent draws of the parameters $\theta_{1}$ and $\theta_{0}$. The upper and lower panels provide the posterior when constraints are and are not accounted for, respectively. The shaded areas represent the feasible regions; points outside the feasible regions are colored in red. This plot is based on 10,000 Gibbs iterations.
@@
Figure 5 demonstrates some of the effects of enforcing these constraints for one particular set [^7] The addition of Laplace noise to the fourth sample moments of ${\mathbf{X}}$ adds additional uncertainty which is not accounted for in the procedure in [5]. Because of this, our rejection sampling approach struggles when the released noisy moments are far from their nonprivate counterparts, leading to the Gibbs sampling procedure being unable to find a draw for the next Gibbs iterate satisfying the constraints and getting “stuck.” For demonstrative purposes, we choose a seed where this issue is not encountered. of released noisy values with $\varepsilon=0.1$ for each of the 11 queries. The shaded triangles in the left panels represent the constraints that $0\leq{\mathbf{Y}}^{\top}{\mathbf{Y}}\leq{\mathbf{Y}}^{\top}{\mathbf{1}}\leq n$. We find that around 20% of the posterior draws without accounting for constraints do not satisfy this inequality, as indicated by the red points. The shaded regions in the right panels represent the feasible region for ${\boldsymbol{\theta}}$. In this case, less than 5% of the posterior draws did not satisfy the constraint in ${\boldsymbol{\theta}}$. But notably, when all relevant constraints are enforced, the posterior distribution is more spread out throughout the feasible region. In this case, not enforcing the constraints leads artificially to an estimated model with too little uncertainty about its parameter estimates and thus underestimated uncertainty around downstream predictions.
@@
This regression example, as well as univariate Gaussian case, suggest that incorporating constraints in Bayesian DP inference, thereby respecting the actual support of the parameters, can result in potentially more accurate representations of posterior distributions. Additionally, the examples indicate that informative prior distributions can have significant impact on DP posterior inferences. The use of weakly informative or default prior distributions may be advantageous, although analysts must take care to ensure posterior distributions are not improper, as demonstrated in Section 4. Finally, these results suggest there is scope for further development of default prior distributions to obtain theoretically valid and accurate posterior distributions with desirable frequentist properties.
Acknowledgments and Disclosure of Funding
@@ Appendix A Re-scaling The Data
We now consider the sensitivities of each of the four statistics. By Lemmas 11 and 12 in [11], the sensitivity of $\bar{Y}$ is $(b-a)/n$ and the sensitivity of $S^{2}$ is $(b-a)^{2}/n$. Similarly, the sensitivity of $\tilde{\bar{Y}}$ is $1/n$ and the sensitivity of $\tilde{S}^{2}$ is $1/n$. Thus, when released under $\varepsilon$-DP via the Laplace Mechanism, the released statistics have distribution
@@ Appendix B The Truncated Gamma Mixture Distribution
Figure 6 compares the probability density of the TGM distribution to the probability density of the gamma distribution with the same $\alpha$ and $\beta$. We see that the shapes are similar, but the rates are different on each side of $\tau$. Notably, the distribution is continuous but is clearly not differentiable at $\tau$.
@@
Algorithm 1 provides a straightforward procedure for sampling from the TGM distribution. It assumes one is able to sample from a truncated gamma distribution, which can be done via standard tools such as the rtrunc function from the truncdist package in R [33].
@@ 10             Sample $X\sim\textsf{Gam}(\alpha,\beta+\lambda)$ truncated to $(\tau,\infty)$
If $\tau\leq 0$, then $X\sim\textsf{Gam}(\alpha,\beta+\lambda)$ and so has PDF
$p(x)=\frac{(\beta+\lambda)^{\alpha}}{\Gamma(\alpha)}x^{\alpha-1}e^{-(\beta+\lambda)x}.$ (14)
Thus, the distribution of $X$ is as desired for $\tau\leq 0$.
@@ Appendix C Derivation of Full Conditionals for Univariate Gaussian Gibbs Sampler
The release mechanism is such that $(\bar{Y}^{*}\perp\!\!\!\perp S^{2*})\mid\bar{Y},S^{2}$. It is convenient to replace the above distribution of $\bar{Y}^{*}$ with the following equivalent formulation, as used in [4, 5, 34].
@@
which we recognize as the full conditional from the public setting, as derived in Chapter 5 of [17]. Similarly, the full conditional for $\sigma^{2}$ is
@@
This is the same full conditional as in the public setting, as derived in Chapter 6 of [17]. Thus, the full conditionals have the form
@@
We recognize this is equivalent to a Gaussian model with known variance and a Gaussian prior on the mean. By Chapter 5 of [17], the full conditional for $\bar{Y}$ is
@@
We recognize this form from the Bayesian LASSO [34], which yields the distribution
@@
If $\varepsilon_{2}>2(n-1)/n$, then the full conditional for $S^{2}$ is not necessarily a TGM without additional assumptions. If reasonable for the application, the analyst could impose the additional constraint that $\sigma^{2}<(n-1)/(n\varepsilon_{2})$, which is sufficient to ensure the full conditional for $S^{2}$ is of the form in (46). Alternatively, if $n$ is large, an analyst could use an approximation of the full conditional for $S^{2}$, such as the proposals of [4, 5].
Appendix D List of Full Conditionals
In this section, we summarize the full conditional under different prior distributions for the DP univariate Gaussian setting with likelihood given by $Y_{i}\overset{iid}{\sim}{\mathcal{N}}_{[0,1]}(\mu,\sigma^{2})$, $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},1/(\varepsilon_{1}n))$, and $S^{2*}\sim\textsf{Lap}(S^{2},1/(\varepsilon_{2}n))$. Subscripts on a distribution denote truncation of the distribution to an interval.

### v2

@@ 5 REGRESSION APPLICATION
The strategies we discuss can be generalized to any Bayesian analysis under DP. To demonstrate, we consider the DP Bayesian linear regression method of Bernstein & Sheldon (2019). In particular, we adapt an example from Bernstein & Sheldon (2019) using data from Brownlee (1965) to illustrate how the strategies for enforcing constraints in Section 3 can be used off-the-shelf to supplement an existing method.
@@
Figure 5: Plot of posterior draws from the linear regression method of Bernstein & Sheldon (2019). The left panels represent draws of the imputed sufficient statistics ${\mathbf{Y}}^{\top}{\mathbf{1}}$ and ${\mathbf{Y}}^{\top}{\mathbf{Y}}$, while the right panels represent draws of the parameters $\theta_{1}$ and $\theta_{0}$. The upper and lower panels provide the posterior when constraints are and are not accounted for, respectively. The shaded areas represent the feasible regions; points outside the feasible regions are colored in red. This plot is based on 10,000 Gibbs iterations.
The data include the response variable cirrhosis rate, $\{y_{i}\}$, explanatory variable drinking rate, $\{x_{i}\}$, and sample size $n=46$. The authors rescale both variables to lie in $[0,1]$ and assume the range is public information. Letting ${\mathbf{Y}}$ be the vector of responses and ${\mathbf{X}}$ be a design matrix with a column of ones, the authors’ model is a simple linear regression of the form ${\mathbf{Y}}\sim{\mathcal{N}}_{n}({\mathbf{X}}{\boldsymbol{\theta}},\sigma^{2}{\mathbf{I}}_{n})$ with the conjugate prior
${\boldsymbol{\theta}}\mid\sigma^{2}\sim{\mathcal{N}}_{2}\left(\begin{bmatrix}1\\ 0\end{bmatrix},\frac{\sigma^{2}}{4}\mathbf{I}_{2}\right),\quad\sigma^{2}\sim\textsf{IG}\left(\frac{40}{2},\frac{1}{2}\right),$ (5)
where $\mathbf{I}_{p}$ is the $p\times p$ identity matrix. Their proposed method involves applying the Laplace Mechanism to each element of the sufficient statistics ${\mathbf{X}}^{\top}{\mathbf{X}}$, ${\mathbf{X}}^{\top}{\mathbf{Y}}$ and ${\mathbf{Y}}^{\top}{\mathbf{Y}}$. To ensure a valid solution exists, they enforce that the matrix ${\mathbf{Z}}^{\top}{\mathbf{Z}}$, where ${\mathbf{Z}}=[{\mathbf{X}}\hskip 5.78172pt{\mathbf{Y}}]$, is positive semi-definite (PSD) by projecting samples to the nearest PSD matrix. They apply their Gibbs-SS-Noisy method, which also adds Laplace noise to the matrix of fourth sample moments of ${\mathbf{X}}$.
Since it is known that $x_{i},y_{i}\in[0,1]$ for all $i$, we can exploit the structure of the linear regression to constrain parameter updates in the Gibbs sampler. In particular, letting ${\mathbf{X}}=[{\mathbf{1}}_{n}\hskip 5.78172pt{\mathbf{x}}_{1}]$, we show in Appendix J that $0\leq{\mathbf{x}}_{1}^{\top}{\mathbf{x}}_{1}\leq{\mathbf{x}}_{1}^{\top}{\mathbf{1}}\leq n$, $0\leq{\mathbf{Y}}^{\top}{\mathbf{Y}}\leq{\mathbf{Y}}^{\top}{\mathbf{1}}\leq n$, ${\mathbf{x}}_{1}^{\top}{\mathbf{Y}}\leq\min\{{\mathbf{x}}_{1}^{\top}{\mathbf{1}},{\mathbf{Y}}^{\top}{\mathbf{1}}\}$, and $40\leq\sigma^{2}\leq 1/$ 4. Letting the regression coefficients be ${\boldsymbol{\theta}}=[\theta_{0}\hskip 5.78172pt\theta_{1}]^{\top}$, we know that $(\theta_{1},\theta_{0})$ must lie in the region depicted in the right panels of Figure 5. We create a constrained Gibbs sampler by drawing each iterate from the unconstrained full conditional distribution and rejecting and resampling whenever all constraints are not satisfied. We also reject and resample if ${\mathbf{Z}}^{\top}{\mathbf{Z}}$ is not PSD.
@@
Figure 5 demonstrates some of the effects of enforcing these constraints for a particular set [^6] The addition of Laplace noise to the fourth sample moments of ${\mathbf{X}}$ adds additional uncertainty which is not accounted for in the procedure in Bernstein & Sheldon (2019). Because of this, our rejection sampling approach struggles when the released noisy moments are far from their nonprivate counterparts, leading to the Gibbs sampling procedure being unable to find a draw for the next Gibbs iterate satisfying the constraints and getting “stuck.” For demonstrative purposes, we choose a seed where this issue is not encountered. of released values with $\varepsilon=0.1$ for each of the 11 queries. The shaded regions in the left panels represent the constraint $0\leq{\mathbf{Y}}^{\top}{\mathbf{Y}}\leq{\mathbf{Y}}^{\top}{\mathbf{1}}\leq n$. As indicated by the red points, nearly 20% of posterior draws without accounting for constraints do not satisfy this inequality. The shaded regions in the right panels represent the feasible region for ${\boldsymbol{\theta}}$. Less than 5% of the posterior draws did not satisfy the constraint in ${\boldsymbol{\theta}}$. But notably, when all relevant constraints are enforced, the posterior distribution is more spread out throughout the feasible region. That is, not enforcing the constraints leads artificially to a model with too little uncertainty about its parameter estimates and thus underestimated uncertainty around downstream predictions.
6 DISCUSSION
A major advantage of the Bayesian paradigm generally is that it facilitates convenient propagation of uncertainty. Indeed, for these reasons, Bayesian inference has been a focus of recent work in the DP literature, e.g., Bernstein & Sheldon (2018), Gong (2022), and Ju et al. (2022). However, inference from these methods may be sensitive to the analyst’s prior selection. As far as we are aware, our work is the first to examine the effect of the prior distribution on inference in settings with substantial noise due to DP and constraints on data values.
@@ 6 DISCUSSION
Our case studies suggest that incorporating constraints in Bayesian DP inference, thereby respecting the actual support of the parameters, can result in potentially more accurate representations of posterior distributions, as well as more accurate point estimates and predictions. The use of weakly informative or default prior distributions may be advantageous, although analysts must ensure posterior distributions are not improper, as demonstrated in Section 4. Finally, these results suggest there is scope for further development of default prior distributions to obtain theoretically valid and accurate posterior distributions with desirable frequentist properties.
In this work, we used the simple setting of bounded data modeled as following a Gaussian distribution as a running case study. Settings where data can be reasonably approximated by a Gaussian distribution are extremely common in applied statistics, but relatively under-explored in the DP literature. This setting also forms the basis for methods applied to a wide array of common tasks in machine learning, such as A/B testing, hierarchical methods, and regression. Additionally, settings where confidential data values are assumed to have an inherent or effective bound are widespread in the DP literature. As far as we are aware, this work is the first to examine the effect of incorporating the bounds as a constraint on a downstream analyst’s model. Extension of the ideas in this work to settings beyond univariate Gaussian and linear regression models is a promising avenue for future work.
Acknowledgements
@@ Appendix A RE-SCALING THE DATA
We now consider the sensitivities of each of the four statistics. By Lemmas 11 and 12 in Du et al. (2020), the sensitivity of $\bar{Y}$ is $(b-a)/n$ and the sensitivity of $S^{2}$ is $(b-a)^{2}/n$. Similarly, the sensitivity of $\tilde{\bar{Y}}$ is $1/n$ and the sensitivity of $\tilde{S}^{2}$ is $1/n$. Thus, when released under $\varepsilon$-DP via the Laplace Mechanism, the released statistics have distribution
@@ Appendix B THE TRUNCATED GAMMA MIXTURE DISTRIBUTION
Figure 6 compares the probability density of a TGM distribution to the probability density of the gamma distribution with the same $\alpha$ and $\beta$. We see that the shapes are similar, but the rates are different on each side of $\tau$. Notably, the distribution is continuous but is clearly not differentiable at $\tau$.
@@
Algorithm 1 provides a straightforward procedure for sampling from the TGM distribution. It assumes one is able to sample from a truncated gamma distribution, which can be done via standard tools such as the rtrunc function from the truncdist package in R (Novomestky & Nadarajah, 2016, License: GPL2absent2$\geq$ 2≥ 2).
@@ 10             Sample $X\sim\textsf{Gam}(\alpha,\beta+\lambda)$ truncated to $(\tau,\infty)$
If $\tau\leq 0$, then $X\sim\textsf{Gam}(\alpha,\beta+\lambda)$, as desired.
@@ Appendix C DERIVATION OF FULL CONDITIONALS FOR UNIVARIATE GAUSSIAN GIBBS SAMPLER
The release mechanism is such that $(\bar{Y}^{*}\perp\!\!\!\perp S^{2*})\mid\bar{Y},S^{2}$. It is convenient to replace the above distribution of $\bar{Y}^{*}$ with the following equivalent formulation, as proposed by Park & Casella (2008) and used in Bernstein & Sheldon (2018, 2019).
@@
which we recognize as the full conditional from the public setting, as derived in Chapter 5 of Hoff (2009). Similarly, the full conditional for $\sigma^{2}$ is
@@
This is the same full conditional as in the public setting, as derived in Chapter 6 of Hoff (2009). Thus, the full conditionals have the form
@@
We recognize this is equivalent to a Gaussian model with known variance and a Gaussian prior on the mean. By Chapter 5 of Hoff (2009), the full conditional for $\bar{Y}$ is
@@
We recognize this form from the Bayesian LASSO (Park & Casella, 2008), which yields the distribution
@@
If $\varepsilon_{2}>2(n-1)/n$, then the full conditional for $S^{2}$ is not necessarily a TGM without additional assumptions. If reasonable for the application, the analyst could impose the additional constraint that $\sigma^{2}<(n-1)/(n\varepsilon_{2})$, which is sufficient to ensure the full conditional for $S^{2}$ is of the form in (45). Alternatively, if $n$ is large, an analyst could use an approximation of the full conditional for $S^{2}$, such as the proposals of Bernstein & Sheldon (2018, 2019).
Appendix D ALGORITHM AND FULL CONDITIONALS
In this section, we provide an explicit algorithm for implementing our Gibbs sampler and enumerate all full conditionals examined in this work. The algorithm is presented below. In the algorithm, we denote the $t^{\textrm{th}}$ draw from the posterior for a variable $x$ as $x_{(t)}$.
Input: Data $(\bar{Y}^{*},S^{2*})$, privacy budget $(\varepsilon_{1},\varepsilon_{2})$, sample size $n$, iterations $T$, hyperparameters $(\mu_{0},\sigma_{0}^{2},\kappa_{0},\nu_{0})$
1 Initialize $\sigma^{2}_{(0)}$, $\bar{Y}_{(0)}$, $S^{2}_{(0)}$, and $\omega^{2}_{(0)}$
2 for $t$ in $1:T$ do
3       Sample $\mu_{(t)}$ from full conditional $(\mu\mid\sigma^{2}=\sigma^{2}_{(t-1)},\omega^{2}=\omega^{2}_{(t-1)},\bar{Y}=\bar{Y}_{(t-1)},S^{2}=S^{2}_{(t-1)},\bar{Y}^{*},S^{2*})$
4       Sample $\sigma^{2}_{(t)}$ from full conditional $(\sigma^{2}\mid\mu=\mu_{(t)},\omega^{2}=\omega^{2}_{(t-1)},\bar{Y}=\bar{Y}_{(t-1)},S^{2}=S^{2}_{(t-1)},\bar{Y}^{*},S^{2*})$
5       Sample $\bar{Y}_{(t)}$ from full conditional $(\bar{Y}\mid\mu=\mu_{(t)},\sigma^{2}=\sigma^{2}_{(t)},\omega^{2}=\omega^{2}_{(t-1)},S^{2}=S^{2}_{(t-1)},\bar{Y}^{*},S^{2*})$
6       Sample $S^{2}_{(t)}$ from full conditional $(S^{2}\mid\mu=\mu_{(t)},\sigma^{2}=\sigma^{2}_{(t)},\omega^{2}=\omega^{2}_{(t-1)},\bar{Y}=\bar{Y}_{(t)},\bar{Y}^{*},S^{2*})$
7       Sample $\omega^{2}_{(t)}$ from full conditional $(\omega^{2}\mid\mu=\mu_{(t)},\sigma^{2}=\sigma^{2}_{(t)},\bar{Y}=\bar{Y}_{(t)},S^{2}=S^{2}_{(t)},\bar{Y}^{*},S^{2*})$
8
9 end for
Output: $(\mu_{(1)},\ldots,\mu_{(T)})$, $(\sigma^{2}_{(1)},\ldots,\sigma^{2}_{(T)})$
Algorithm 2 Run our proposed Gibbs sampler.
In Algorithm 2, the first step is to initialize values for each of the variables, which we describe further in the following paragraph. The analyst then repeatedly samples from the full conditionals for each of the variables. The forms of the full conditionals under different assumptions are described below. The final step is to output the posterior draws for the two parameters.
The posterior distribution is not very sensitive to the initial values, but well-thought-out values may reduce the runtime. In our examples, we set $\sigma^{2}_{(0)}=S^{2}_{(0)}=S^{2*}$ if $S^{2*}\in(0,1/4)$ and $\bar{Y}_{(0)}=\bar{Y}^{*}$ if $\bar{Y}^{*}\in[0,1]$. If $\bar{Y}^{*}$ or $S^{2*}$ are not in these regions, they are set near the closest boundary (although note that $\sigma^{2}_{(0)}$ and $S^{2}_{(0)}$ should not be set exactly equal to zero). We set $\omega^{2}_{(0)}=2/(\varepsilon_{1}^{2}n^{2})$.
We now summarize the full conditional under different prior distributions for the DP univariate Gaussian setting with likelihood given by $Y_{i}\overset{iid}{\sim}{\mathcal{N}}_{[0,1]}(\mu,\sigma^{2})$, $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},1/(\varepsilon_{1}n))$, and $S^{2*}\sim\textsf{Lap}(S^{2},1/(\varepsilon_{2}n))$. Subscripts on a distribution denote truncation of the distribution to an interval.